CS6380 Lab Report A1 Your Name Date

1 Introduction

Briefly summarize the nature of the system, the basic numerical or other algorithm, and the interesting or relevant questions.

Reference Turing[1], and for figures, see Figure 1.

Figure 1: Linear Set of Cells.

2 Method

Describe the algorithm and how it is implemented in the program, but do not give the program itself. Explain any special features of the implementation. The program should follow the coding style of the class:

• Only use Matlab functions (i.e., no scripts)

• Unless the function requires it,

– do not: Write to the interpreter

– do not: Read or write files

– do not: Draw, plot, etc.

• do not: exceed 72 characters on a source line

• do: use required header, and indent it correctly

3 Verification of Program

Confirm that your program is not incorrect by considering special cases and by giving at least one comparison to a hand calculation or known result.

4 Data

Show the results of some typical runs in graphical or tabular form. Additional runs can be included in an appendix. All runs should be labeled, and all tables and figures must be referred to in the body of the text. Each figure and table should have a caption with complete information, e.g., the value of the set parameters, and units on the axes.

5 Analysis

In general, the analysis of your results will include a determination of qualitative and quantitative relationships between variables, and any necessary numerical estimates.

6 Interpretation

Summarize your results and explain them in simple physical terms whenever possible. Specific questions that were raised in the assignment should be addressed here. Also, give suggestions for future work or possible extensions.

7 Critique

Summarize the important concepts related to the problem or method for which you gained a better understanding and discuss the specific techniques that you learned. Make specific comments on the problem and your suggestions for improvements or alternatives to your approach.

8 Log

Keep a log of the time spent on each problem and include it in the report.

References

[1] Alan Turing. The Chemical Basis of Morphogenesis. Philosophical Transactions of the Royal Society of London, Series B, Vol. 237(641):37–72, 1952.